Out of 450 people sampled, 113 had kids. Based on this, construct a 99% confidence interval for the true population proportion of people with kids.
Answer #1

We have given here,                          
                          
x=113                      
n=450                      
Estimate for sample proportion                          
\hat{p}=\frac{x}{n}=\frac{ 113 }{ 450 }= 0.2511
Level of significance is =1-0.99=0.01                  
Z critical value(using Z table)=2.576                  

\hat{p}\pm Z*\sqrt{\frac{\hat{p}*(1-\hat{p})}{n}}= 0.2511 \pm 2.576 \sqrt{\frac{\ 0.2511 * 0.7489 }{ 450 }}=( 0.1984 , 0.3038 )